# Conceiving Naturally After IVF:

### v1

@@ Conclusion
We find that conceiving a baby via ART does not affect the risk of preterm birth, preterm spontaneous labour, the gestational age at birth, the birth weight of the baby nor their APGAR scores, in subfertile mothers. Our estimated null effects are precise, leaving us unable to rule out only ART effects that are clinically small 121212 We can exclude the following effect sizes for the effect of ART on obstretric outcomes: effects between [-0.16, 0.10] weeks for gestational age at birth; between [-1.1, 1.9] p.p. for the risk of preterm birth; between [-0.07, 0.09] points for the APGAR1 score; between [-0.02, 0.1] points for the APGAR5 score; between [-14, 60] grams for the effect of ART on birth weight. . Moreover, we find that ART has a small negative impact (${}-5p.p.$) on the chance of a c-section and of an induction of labour. While the sign of these effect is robust to large levels of unobserved confounding, such large levels of confounding effects would further reduce the magnitude of these estimates and make them statistically insignificant.

### v2

@@ Conclusion
We find that conceiving a baby via ART does not affect the risk of preterm birth, preterm spontaneous labour, the gestational age at birth, the birth weight of the baby nor their APGAR scores, in subfertile mothers. Our estimated null effects are precise, leaving us unable to rule out only ART effects that are clinically small 131313 We can exclude the following effect sizes for the effect of ART on obstretric outcomes: effects between [-0.16, 0.10] weeks for gestational age at birth; between [-1.1, 1.9] p.p. for the risk of preterm birth; between [-0.07, 0.09] points for the APGAR1 score; between [-0.02, 0.1] points for the APGAR5 score; between [-14, 60] grams for the effect of ART on birth weight. . Moreover, we find that ART has a small negative impact (${}-5p.p.$) on the chance of a c-section and of an induction of labour. While the sign of these effect is robust to large levels of unobserved confounding, such large levels of confounding effects would further reduce the magnitude of these estimates and make them statistically insignificant.